Both cavities are locked at the same time. The temperature setting are, RCAV = 34.95, ACAV = 37.2.
I realigned the beam onto the PD to get maximum contrast. I'll readjust the setting back to the original value
and see if the beat noise is improved.
I just notice that one of the beam on the mirror on ACAV's path behind the cavity is almost clipped. I'll readjust it tomorrow.
Beat measurement optics' height is changed to 3". I cleaned all optics already, but I couldn't really clean 1/2 and 1/4 wave plates, one of the f =200 mm lens is quite hard to clean.
I'll wait and ask someone before trying to clean again. I cannot lock both cavities at the same time, once I can, I'll align the beam on the PD.
Also ACAV's PD for ACAV_trans_PD is broken. It gives out 11 V regardless of the beam falling on the PD, so I replace it with a PD that is used for NPRO_PWRMON.
I designed the layout for optics behind the cavities for beat measurement, and calculated the mode matching for the beam.
Since the current optics height for beat is quite high (7 inches), we want to lower it to 3 inches, make it more symmetric, and more compact.
The PD's diameter is 300 mm, so the beam spot on it will be ~50um.
All the lenses I need are prepared.